# Can Large Language Models Infer Causation from Correlation?

### v2

@@ Abstract
Causal inference is one of the hallmarks of human intelligence. While the field of CausalNLP has attracted much interest in the recent years, existing causal inference datasets in NLP primarily rely on discovering causality from empirical knowledge (e.g., commonsense knowledge). In this work, we propose the first benchmark dataset to test the pure causal inference skills of large language models (LLMs). Specifically, we formulate a novel task Corr2Cause, which takes a set of correlational statements and determines the causal relationship between the variables. We curate a large-scale dataset of more than 200K samples, on which we evaluate seventeen existing LLMs. Through our experiments, we identify a key shortcoming of LLMs in terms of their causal inference skills, and show that these models achieve almost close to random performance on the task. This shortcoming is somewhat mitigated when we try to re-purpose LLMs for this skill via finetuning, but we find that these models still fail to generalize – they can only perform causal inference in in-distribution settings when variable names and textual expressions used in the queries are similar to those in the training set, but fail in out-of-distribution settings generated by perturbing these queries. Corr2Cause is a challenging task for LLMs, and would be helpful in guiding future research on improving LLMs’ pure reasoning skills and generalizability.111 Our data is at https://huggingface.co/datasets/causalnlp/corr2cause.
@@ 1 Introduction
Based on our Corr2Cause dataset with 200K samples, we investigate two main research questions: (1) How well do existing LLMs perform on this task? (2) Can existing LLMs be re-trained or re-purposed on this task and obtain robust causal inference skills? Through extensive experiments, we show empirically that none of the seventeen existing LLMs we investigate perform well on this pure causal inference task. We also show that although LLMs can demonstrate better performance after being finetuned on the data, the causal inference skills attained by them are not robust. In summary, our contributions are as follows:
@@
    We evaluate the performance of seventeen LLMs on our dataset, finding that all of them perform poorly, close to the random baseline;
@@ 3.7 Statistics of the Resulting Data
We show the statistics of our Corr2Cause dataset in Table 3. Overall, our dataset contains 207,972 samples, where 18.57% of the samples have the positive label (i.e., with validity=1). The average length of the premise is 424.11 tokens, and hypothesis 10.83 tokens. We split the data into 205,734 training samples, 1,076 development and 1,162 test samples. Since the main purpose of the dataset is to benchmark the performance of LLMs, we prioritize the test and development sets to have a comprehensive coverage over all sizes of graphs. Specifically, we iterate through the subset of our data for each $N$, and split it entirely for only the test and development sets if the data is less than 1K, which is the case for $N=2$ and $3$. For the other subsets that are larger, we randomly sample up to 1K or 10% of the data, whichever is smaller, to the test and development sets. We set the cap to be 1K in order to form a reasonable computation budget, since many LLMs are expensive to query in the inference mode. Aside from the test and valid sets, all the rest of the data goes into the training set.
@@
Note for our dataset v2.0: We notice that our original dataset (v1.0) has duplication due to symmetric relations and verbalizations of the hypothesis. E.g., Is-Parent(A, B) has the exact hypothesis verbalization as Is-Child(B, A). Hence, for our dataset v2.0, we perform a careful de-duplication, and update the data statistics in Table 3. See more version comparison details in Appendix D. Note that, due to the symmetry, the current version is a random sample half of the size of the original version, so the modeling results in the experiment section roughly hold.
@@ 4.1 Experimental Setup
We set up a diverse list of LLMs for the experiments on our Corr2Cause dataset. To test existing LLMs, we first include six commonly used BERT-based NLI models in the transformers library (Wolf et al., 2020) with the most number of downloads: BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019), BART (Lewis et al., 2020), DeBERTa (He et al., 2021), DistilBERT (Sanh et al., 2019), and DistilBART (Shleifer & Rush, 2020). Apart from these BERT-based NLI models, we also evaluate the general-purpose autoregressive LLMs based on GPT (Radford et al., 2019): GPT-3 Ada, Babbage, Curie, Davinci (Brown et al., 2020); its instruction-tuned versions (Ouyang et al., 2022), text-davinci-001, text-davinci-002, and text-davinci-003; and GPT-3.5 (i.e., ChatGPT), and the latest GPT-4 (OpenAI, 2023) by April 2023, using the OpenAI API [^2] https://openai.com/api/ with temperature 0. We also evaluate the recent, more efficient models, LLaMa (Touvron et al., 2023) and Alpaca (Taori et al., 2023).
When inspecting the behavior of finetuned models, we adopt a large set of models, including GPT-based models (GPT-3 Ada, Babbage, Curie, and Davinci) using the OpenAI finetuning API for classification, [^3] https://platform.openai.com/docs/guides/fine-tuning open-sourced decoder-only models (GPT2, GPT2-Large, GPT2-XL, LLaMA-7B, and LLaMa2-7B), BERT-based models from scratch (BERT-Base, BERT-Large, RoBERTa-Base, and RoBERTa-Large), and BERT-Based NLI models (BERT-Base MNLI, BERT-Large MNLI, RoBERTa-Base MNLI, and RoBERTa-Large MNLI) using the transformers library (Wolf et al., 2020). Our training details are available in Appendix A.
For the random baselines, we provide “always majority” to predict the majority class 100% of the time, “random (uniform)” which randomly samples a label with 50% chance for each, and “random (proportional)” which samples a label from a Bernouli distribution proportional to the development set label distribution.
@@ 4.2 The Corr2Cause Skill in Existing LLMs
We show the performance of 17 LLMs in Table 4. We can see that pure causal inference is a very challenging task across all existing LLMs. Among all the LLMs, the best performance is 33.38% F1 by BART MNLI, which is even higher than latest GPT-based model, GPT-4. Notably, many models are worse than random guess, which means that they totally fail at this pure causal inference task.
@@ 4.3 Finetuned Performance
Next, we address the question: Can we re-purpose LLMs to learn this task?
The experimental results in Table 4(a) of 17 models finetuned on our Corr2Cause seem very strong at first sight. Most models see a substantial increase, among which the finetuned BERT-based NLI models demonstrate the strongest performance. The best-performing one, RoBERTa-Large MNLI, achieves 94.74% F1 score on this task, as well as very high precision, recall and accuracy scores.
@@ 4.4 Fine-Grained Performance by Causal Relation
In addition to the overall results mentioned above, we also conduct a fine-grained analyze to check the performance of the strongest model, RoBERTa-Large MNLI, by our six causal relation types. As in Table 5(a), the model is very good at judging relations such as Is-Parent, Is-Descendant and Has-Confounder, all with more than 96% F1 scores, whereas it is several points weaker on the Has-Collider relations. This could be due to that the collider relation is the most special type, requiring identification of the V-structure based on both the unconditional independence based on the two variables only and correlations whenever conditioned on a common descendant.
@@ 4.5 Robustness Analysis
Specifically, we adopt the common setup of text adversarial attack (Morris et al., 2020; Jin et al., 2020) to preserve the training set and keep the same saved models, but run the inference n the perturbed test set. In this way, we separate the possibilities of the models only overfitting on the training data vs. mastering the reasoning skills.
@@
We can see from Table 4(b) that all the models drop drastically, by up to 39.29 when we paraphrase the test set, and they decrease substantially by up to 62.3 when we refactor the variable names. The best-performing model, RoBERTa-Large MNLI, is especially sensitive towards paraphrasing, demonstrating the most drop among all models; however, it is the most robust against the variable refactorization, maintaining a high F1 score of 67.87. We conduct fine-grained analysis for RoBERTa-Large MNLI under perturbation in Table 5(b). We can see the the main source of the performance drop of the model comes from the two classes, Is-Ancestor (decreasing to 45.45%) and Is-Descendant (decreasing to 29.41%), while the other classes stay relatively robust, keeping their F1 scores over 70%.
From this analysis, we make the following suggestions to future studies testing this Corr2Cause skill of LLMs. First, it is safe to use it as a test set to benchmark existing LLMs’ performance, since the data we generate is out-of-distribution from the training data of the current LLMs. Then, when testing finetuned models, it is very important to accompany adversarial attack together with the i.i.d. test set. We also provide our perturbed versions of the test set in our data for future work to test the generalizability skill.
@@ 5 Related Work
Another related area of literature is logical and inference tasks. A well-established task is natural language inference (NLI), which identifies the semantic relationship between a pair of sentences (MacCartney & Manning, 2008; Bowman et al., 2015). NLI datasets mainly focus on the set and paraphrase relations, such as “a group of boys are playing football” can entail “some guys are playing football,” where “boys” are a sub-concept of “guys” and “a group of” and “some” are paraphrases. Existing datasets cover entailment in news articles (Dagan et al., 2006), image captions (Bowman et al., 2015), and across multiple genres (Williams et al., 2018). Recently, there has been increasing efforts to extend the inference task to various logical inference skills such as deductive logic and propaganda techniques (Jin et al., 2022; Alhindi et al., 2022). Our Corr2Cause dataset is the first dataset testing the correlation-to-causation inference skill, which is unique of its type.
6 Limitations and Future Work
@@
7 Conclusion
In this work, we introduced a novel task, Corr2Cause, to infer causation from correlation, and collected a large-scale dataset of over 200K samples. We evaluated an extensive list of LLMs on this new task, and showed that off-the-shelf LLMs perform poorly on this task. We also show that it is possible to re-purpose LLMs on this task by finetuning, but future work needs to be aware of the out-of-distribution generalization problem. To avoid the Goodhart’s law, we recommend using this dataset to benchmark the pure causal inference skills for LLMs that have not seen this dataset. Given the limited reasoning abilities of current LLMs, and the difficulty of separating actual reasoning from training-corpus-derived knowledge, it is imperative that our community focus on work aiming to accurately disentangle and measure both abilities. We believe that the present work is a first such step.
@@ Appendix B Generating Natural Stories
Example 1 (Label=Negative)
Example 2 (Label=Positive)
Symbolic Form
Premise: Suppose there is a closed system of 2 variables, A and B. All the statistical relations among these 2 variables are as follows: A correlates with B.
@@
Natural Story
Premise: Suppose there is a closed system of 2 variables, ice cream sales and swimming pool attendance. All the statistical relations among these 2 variables are as follows: ice cream sales correlate with swimming pool attendance.
@@ Appendix D Change Log for the Dataset Version Update
Two Equivalent Forms
Duplication Property
De-Duplication Method
{$\ \end{array}\right.$ Is-Parent(i, j) Two exact same strings Keep only one, by forcing $i<j$
@@
{$\ \end{array}\right.$ Is-Ancestor(i, j) (Original)
Two different strings, but
Randomly sample one out of the two
Is-Descendent(j, i) (Original)
semantically equivalent
{$\ \end{array}\right.$ Is-Ancestor(i, j) (Paraphrased) Two exact same strings Keep only one, by forcing $i<j$
@@
{$\ \end{array}\right.$ Has-Collider(i, j)
Two different strings, but
Randomly sample one out of the two
Has-Collider(j, i)
semantically equivalent
{$\ \end{array}\right.$ Has-Confounder(i, j)
Two different strings, but
Randomly sample one out of the two
Has-Confounder(j, i)
semantically equivalent

### v3

@@ Abstract
Causal inference is one of the hallmarks of human intelligence. While the field of Causal NLP has attracted much interest in the recent years, existing causal inference datasets in NLP primarily rely on discovering causality from empirical knowledge (e.g., commonsense knowledge). In this work, we propose the first benchmark dataset to test the pure causal inference skills of large language models (LLMs). Specifically, we formulate a novel task Corr2Cause, which takes a set of correlational statements and determines the causal relationship between the variables. We curate a large-scale dataset of more than 200K samples, on which we evaluate seventeen existing LLMs. Through our experiments, we identify a key shortcoming of LLMs in terms of their causal inference skills, and show that these models achieve almost close to random performance on the task. This shortcoming is somewhat mitigated when we try to re-purpose LLMs for this skill via finetuning, but we find that these models still fail to generalize – they can only perform causal inference in in-distribution settings when variable names and textual expressions used in the queries are similar to those in the training set, but fail in out-of-distribution settings generated by perturbing these queries. Corr2Cause is a challenging task for LLMs, and can be helpful in guiding future research on improving LLMs’ pure reasoning skills and generalizability.111 Our data is at https://huggingface.co/datasets/causalnlp/corr2cause.
@@ 1 Introduction
Based on our Corr2Cause dataset with 200K samples, we investigate two main research questions: (1) How well do existing LLMs perform on this task? (2) Can existing LLMs be re-trained or re-purposed on this task and obtain robust causal inference skills? Through extensive experiments, we show empirically that none of the 17 existing LLMs we investigate perform well on this pure causal inference task. We also show that although LLMs can demonstrate better performance after being finetuned on the data, the causal inference skills attained by them are not robust. In summary, our contributions are as follows:
@@
    We evaluate the performance of 17 LLMs on our dataset, finding that all of them perform poorly, close to the random baseline;
@@ 3.7 Statistics of the Resulting Data
We show the statistics of our Corr2Cause dataset in Table 3. Overall, our dataset contains 207,972 samples, where 18.57% of the samples have the positive label (i.e., with validity=1). The average length of the premise is 424.11 tokens, and hypothesis 10.83 tokens. We split the data into 205,734 training samples, 1,076 development and 1,162 test samples. [^2] Note for our dataset v2.0: We notice that our original data (v1.0) has duplication due to symmetric relations and verbalizations of the hypothesis. E.g., Is-Parent(A, B) has the exact hypothesis verbalization as Is-Child(B, A). Hence, for our data v2.0, we perform a careful de-duplication, and update the data statistics in Table 3. See more version comparison details in Appendix D. Note that, due to the symmetry, the current version is a random sample half of the size of the original version, so the modeling results in the experiment section roughly hold. Since the main purpose of the dataset is to benchmark the performance of LLMs, we prioritize the test and development sets to have a comprehensive coverage over all sizes of graphs. Specifically, we iterate through the subset of our data for each $N$, and split it entirely for only the test and development sets if the data is less than 1K, which is the case for $N=2$ and $3$. For the other subsets that are larger, we randomly sample up to 1K or 10% of the data, whichever is smaller, to the test and development sets. We set the cap to be 1K in order to form a reasonable computation budget, since many LLMs are expensive to query in the inference mode. Aside from the test and valid sets, all the rest of the data goes into the training set.
@@ 4.1 Experimental Setup
We set up a diverse list of LLMs for the experiments on our Corr2Cause dataset. To test existing LLMs, we first include six commonly used BERT-based NLI models in the transformers library (Wolf et al., 2020): BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019), BART (Lewis et al., 2020), DeBERTa (He et al., 2021), DistilBERT (Sanh et al., 2019), and DistilBART (Shleifer & Rush, 2020). Apart from these BERT-based NLI models, we also evaluate the general-purpose autoregressive LLMs based on GPT (Radford et al., 2019): GPT-3 Ada, Babbage, Curie, Davinci (Brown et al., 2020); its instruction-tuned versions (Ouyang et al., 2022), text-davinci-001, text-davinci-002, and text-davinci-003; and GPT-3.5 (i.e., ChatGPT), and the latest GPT-4 (OpenAI, 2023) by April 2023, using the OpenAI API (https://openai.com/api/) with temperature 0. We also evaluate the recent, more efficient models, LLaMa (Touvron et al., 2023) and Alpaca (Taori et al., 2023).
When inspecting the behavior of finetuned models, we adopt a large set of models, including GPT-based models (GPT-3 Ada, Babbage, Curie, and Davinci) using the OpenAI finetuning API for classification at https://platform.openai.com/docs/guides/fine-tuning, open-sourced decoder-only models (GPT2, GPT2-Large, GPT2-XL, LLaMA-7B, and LLaMA2-7B), BERT-based models from scratch (BERT-Base, BERT-Large, RoBERTa-Base, and RoBERTa-Large), and BERT-Based NLI models (BERT-Base MNLI, BERT-Large MNLI, RoBERTa-Base MNLI, and RoBERTa-Large MNLI) using the transformers library (Wolf et al., 2020). See training details in Appendix A.
For the random baselines, we provide “always majority” to predict the majority class 100% of the time, “random (uniform)” to uniformly sample a label (i.e., 50% for each), and “random (proportional)” to sample a label from a Bernouli distribution proportional to the development set label distribution.
@@ 4.2 The Corr2Cause Skill in Existing LLMs
We show the performance of seventeen LLMs in Table 4. We can see that pure causal inference is a very challenging task across all existing LLMs. Among all the LLMs, the best performance is 33.38% F1 by BART MNLI, which is even higher than the latest GPT-based model, GPT-4. Notably, many models are worse than random guess, which means that they totally fail at this pure causal inference task. The observation still holds for few-shot chain-of-thought prompts tested in Appendix G.
@@ 4.3 Finetuned Performance
Next, we address the question: Can we re-purpose LLMs to learn this task? The experimental results in Table 5(a) of 17 models finetuned on our Corr2Cause seem very strong at first sight. Most models see a substantial increase, among which the finetuned BERT-based NLI models demonstrate the strongest performance. The best-performing one, RoBERTa-Large MNLI, achieves 94.74% F1 score on this task, as well as very high precision, recall and accuracy scores.
@@ 4.4 Fine-Grained Performance by Causal Relation
In addition to the overall results mentioned above, we conduct a fine-grained analysis to check the performance of the strongest finetuned model, RoBERTa-Large MNLI, by our six causal relation types. As in Table 6(a), the model is very good at judging relations such as Is-Parent, Is-Descendant and Has-Confounder, all with more than 96% F1 scores, whereas it is several points weaker on the Has-Collider relations. This could be due to that the collider relation is the most special type, requiring identification of the V-structure based on both the unconditional independence based on the two variables only and correlations whenever conditioned on a common descendant. We also conduct error analysis for non-finetuned models in Appendix F.
@@ 4.5 Robustness Analysis
Specifically, we adopt the common setup of text adversarial attack (Morris et al., 2020; Jin et al., 2020) to preserve the training set and keep the same saved models, but run the inference on the perturbed test set. In this way, we separate the possibilities of the models only overfitting on the training data vs. mastering the reasoning skills.
@@
We can see from Table 5(b) that all the models drop drastically, by up to 39.29 on the paraphrased test set, and up to 62.30 after variable refactorization. The best-performing model, RoBERTa-Large MNLI, is especially sensitive towards paraphrasing, demonstrating the most drop among all models; however, it is the most robust against the variable refactorization, maintaining a high F1 score of 67.87. We conduct fine-grained analysis for RoBERTa-Large MNLI under perturbation in Table 6(b). We can see the the main source of the performance drop of the model comes from the two classes, Is-Ancestor (decreasing to 45.45%) and Is-Descendant (decreasing to 29.41%), while the other classes stay relatively robust, keeping their F1 scores over 70%.
From this analysis, we make the following suggestions to future studies testing this Corr2Cause skill of LLMs. First, it is safe to use it as a test set to benchmark existing LLMs’ performance, since the data we generate is out-of-distribution from the training data of the current LLMs. Then, when testing finetuned models, it is very important to accompany adversarial attack together with the i.i.d. test set. We open-source our perturbed test sets for future work to test the generalizability skill.
@@ 5 Related Work
Another related area of literature is logical and inference tasks, of which a well-established one is natural language inference (NLI), to identify the semantic relationship between a pair of sentences (MacCartney & Manning, 2008; Bowman et al., 2015). NLI datasets mainly focus on the set and paraphrase relations. For example, “a group of boys are playing football” can entail “some guys are playing football,” where “boys” are a sub-concept of “guys,” and “a group of” and “some” are paraphrases. Recently, there have been increasing efforts to extend the inference task to various logical inference skills such as deductive logic and propaganda techniques (Jin et al., 2022; Alhindi et al., 2022). Our Corr2Cause dataset is the first dataset testing the correlation-to-causation inference skill, which is unique of its type.
6 Conclusion
In this work, we introduced a novel task, Corr2Cause, to infer causation from correlation, and collected a large-scale dataset of over 200K samples. We evaluated an extensive list of LLMs on this new task, and showed that off-the-shelf LLMs perform poorly on this task. We also show that it is possible to re-purpose LLMs on this task by finetuning, but future work needs to be aware of the out-of-distribution generalization problem. To avoid the Goodhart’s law, we recommend using this dataset to benchmark the pure causal inference skills for LLMs that have not seen this dataset. Given the limited reasoning abilities of current LLMs, and the difficulty of separating actual reasoning from training-corpus-derived knowledge, it is imperative that our community focus on work aiming to accurately disentangle and measure both abilities. We believe the present work is a first such step.
Limitations and Future Work
@@ 6 Conclusion
Acknowledgment
We thank Riley Goodside for valuable suggestions to improve our prompts to LLMs. We thank Luigi Gresele and Amir Hossein Karimi for their suggestions to help us improve the formulation of our causal discovery questions.
This material is based in part upon work supported by the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039B; by the Machine Learning Cluster of Excellence, EXC number 2064/1 – Project number 390727645; by a National Science Foundation award (#2306372); by a Swiss National Science Foundation award (#201009) and a Responsible AI grant by the Haslerstiftung. Zhijing Jin is supported by PhD fellowships from the Future of Life Institute and Open Philanthropy. We also thank OpenAI for granting Zhijing quota to their API of GPT series through the Researcher Access Program.
@@ Appendix B Generating Natural Stories
Example 1 (Label=Negative) Example 2 (Label=Positive)
Symbolic Form Premise: Suppose there is a closed system of 2 variables, A and B. All the statistical relations among these 2 variables are as follows: A correlates with B.
@@
Natural Story Premise: Suppose there is a closed system of 2 variables, ice cream sales and swimming pool attendance. All the statistical relations among these 2 variables are as follows: ice cream sales correlate with swimming pool attendance.
@@ Appendix D Change Log for the Dataset Version Update
Two Equivalent Forms Duplication Property De-Duplication Method
{$\ \vphantom{1}\end{array}\right.$ Is-Parent(i, j) Two exact same strings Keep only one, by forcing $i<j$
@@
{$\ \vphantom{1}\end{array}\right.$ Is-Ancestor(i, j) (Original) Two different strings, but Randomly sample one out of the two
Is-Descendent(j, i) (Original) semantically equivalent
{$\ \vphantom{1}\end{array}\right.$ Is-Ancestor(i, j) (Paraphrased) Two exact same strings Keep only one, by forcing $i<j$
@@
{$\ \vphantom{1}\end{array}\right.$ Has-Collider(i, j) Two different strings, but Randomly sample one out of the two
Has-Collider(j, i) semantically equivalent
{$\ \vphantom{1}\end{array}\right.$ Has-Confounder(i, j) Two different strings, but Randomly sample one out of the two
Has-Confounder(j, i) semantically equivalent
@@ Appendix F Fine-Grained Error Analysis
Appendix G LLM Performance Optimization
Since our experiments in Section 4.2 are based on plain, zero-shot prompts, we explore whether better prompting strategies could improve the performance. We enhance the query prompt by incorporating several strategies: (1) Utilizing a system prompt that specifies the model’s expertise (“You are a highly intelligent question-answering bot with profound knowledge of causal inference.”); (2) Including a pair of few-shot examples, one positive and one negative; (3) Implementing chain-of-thought prompting with “Let’s think step by step.” to encourage the language model to generate step-by-step reasoning. In Table 13, we present the evaluation results on the relatively affordable model, GPT-3.5, where the optimized prompt leads to a 4-point improvement in F1 over the original performance. However, we can see that despite the deployment of all three strategies, the model continues to struggle with this challenging task.
F1 Precision Recall Accuracy
GPT-3.5 (plain query; original) 21.69 17.79 27.78 69.46
GPT-3.5 (enhanced query) 25.44 17.29 48.11 52.01
Table 13: Performance of GPT-3.5 with different queries. We quote the original performance from Table 4.